Name: ID: A In summer squash, there are two genes that determine fruit color. The two genes sort independently. Two white-fruited plants are crossed. Both parents are known to be heterozygous for both genes The cross produces the following offispring: 20 green-fruited plants, 58 yellow-fruited plants, and 218 white-fruited plants. 13. A. (3pts) Using the letters A/a and B/b to indicate the two different genes, draw a Punnett square for the cross of the two white-fruited plants B. (4pts) List the genotype classes in the offispring AND give the corresponding phenotype of each. C. (1pt) Name the kind of epistastic interaction (of the 4 you are required to know) that best explains the phenotypic ratios you obtained D. (3pts) Perform a chi-square goodness of fit test to determine if the phenotypic frequencies you obtained fit the epistasis interaction you chose. Show your work E. (2pts) State whether you reject or fail to reject the null (chance) hypothesis
Answer #1

A)      Gene pair 1: White dominant =A, coloured recessive=a

        Gene pair 2: Yellow dominant =B, green recessive=b

AaBb X AaBb White White AB AB AABB Ab AABb Ab AaBIB AaBb Ab AABb AAbb AaBb Aabb AaBB AABb aaBB aaBb AaBb Aabb aaBb aabb

B)     

AABB:White

AABb : White

AaBB: White

AaBb: White

AAbb: White

Aabb: White

aaBB:Yellow

aaBb: Yellow

aabb: Green

C)      Epistatic interaction type: Complete dominance at both gene pairs, but one gene, when dominant, epistatic to the other.

Phenotypic ratio Expected white:yellow:green = 12:3:1

Phenotypic ratio observed 218:58:20 ῀ 12:3:1

D)      The expected number for different phenotype is

White= (296/16)*12= 222

Yellow= (277/16)*3= 55

green= (277/16)*1=19

Observed

Expected

Row Totals

White

218

222

440

Yellow

58

55

113

Green

20

19

39

Column Totals

296

296

592  (Grand Total)

The chi-square is 0.1417.

p-value is 0.93. The result is not significant at p< 0.05

E)      Thus, null hypothesis ‘both sets of data are different’ rejected and alternative hypothesis ‘both sets of data are not different’ accepted.